Which of the following statements about Hardy Weinberg equilibrium is true?

a.

It explains why dominant alleles do not necessarily replace recessive alleles in a population.

b.

It applies only to populations in which there is gene flow.

c.

It assumes that populations are small.

d.

It assumes that individuals prefer to mate with individuals with certain genotypes.

e.

None of the above.

An answer and explanation would be greatly appreciated, thank you!
Answer #1

Answer:

d. It assumes that individuals prefer to mate with individuals with certain genotypes.

Explanation:

H-W equlibrium = p+q=1

p = Dominant allele frequency

q = recessive allele frequency

The over all allele's frequency is equal to one. It means that that inviduals prefer to mate with individuals with certain genotypes.